OpenStudy (anonymous):

Please Help me.. :) (2/x^28x+15) +(1/x^2+11x+30)

OpenStudy (saifoo.khan):

Factor the denominators and then find the LCD.

OpenStudy (anonymous):

They don't have a LCD because some of the numbers are prime.

OpenStudy (saifoo.khan):

none of them is prime

OpenStudy (anonymous):

You said i should find the LCD but 11x is only divisible by it self.

OpenStudy (saifoo.khan):

\[\frac{2}{(x+3)(x+5)} + \frac{1}{(x+5)(x+6)}\]Can you solve now?

OpenStudy (anonymous):

I can try. Okay.

OpenStudy (anonymous):

I took the function you gave me and i took the 2 and divided it into the six leaving them to be the same. So do i go ahead and add them straight across

OpenStudy (saifoo.khan):

o_O

OpenStudy (anonymous):

Is that correct? I got 2/(x+5) (x+3)

OpenStudy (saifoo.khan):

no

OpenStudy (anonymous):

2/(x+10) (x+9)

OpenStudy (saifoo.khan):

Nope

OpenStudy (anonymous):

What am i doing wrong?

OpenStudy (anonymous):

The function you gave me should i have just continued to add straight across?

OpenStudy (saifoo.khan):

Yes, but u have to take the LCD first

OpenStudy (anonymous):

Is the LCD is (x+5)

OpenStudy (saifoo.khan):

(x+5)(x+3)(x+6)

OpenStudy (anonymous):

3/(x+5)(x+3)(x+6)

OpenStudy (saifoo.khan):

\[\frac{2(x+6) + 1(x+3)}{(x+5)(x+3)(x+6)}\]

OpenStudy (anonymous):

Now do i cancel them or multiply?

OpenStudy (saifoo.khan):

Just simplify.

OpenStudy (anonymous):

I got two different answer: The first one i cancel them out i got 3/(x+5) The second answer i multiplied straight across then cancel out (x+2)/(x+5) Is any of them correct?

OpenStudy (saifoo.khan):

none

OpenStudy (saifoo.khan):

\[\frac{2x+12+ x + 3}{(x+5)(x+3)(x+6)}\] \[\frac{3x+15}{(x+5)(x+3)(x+6)}\] \[\frac{3(x+5)}{(x+5)(x+3)(x+6)}\]

OpenStudy (saifoo.khan):

\[\frac{3}{(x+3)(x+6)}\]

OpenStudy (anonymous):

I saw how you got it but that's what i got too.

OpenStudy (anonymous):

Thank you for all your help. I really appreciated it.

OpenStudy (saifoo.khan):

Welcome. :)